A current-carrying conductor made of aluminum grad
Answer #1


current density J = I/A

in region 1


J1 = I/A1

A1 = I/J1


dirft speed v1 = I/(n*e*A1)

v1 = I*J1/(n*e*I) = J1/(n*e)


I = current in the conductor

n = electron density = 6*10^28

v1 = 3.8*10^3/(6*10^28*1.6*10^-19)


v1 = 3.95*10^-7 m/s


in region 2


A2 = A1/2 = I/(2*J1)

v2 = 2*J2/(n*e) = 2*3.8*10^3/(6*10^28*1.6*10^-19)


v2 = 7.92*10^-7 m/s